OpenStudy (anonymous):

A company sells boxes of small blue pom-poms (b) for $42 and boxes of large orange pom-poms (p) for $52. They make batches of small blue pom-poms that fill 8 boxes and batches of large orange pom-poms that fill 4 boxes. The company only has 40 boxes. Which system of equations helps the company plan to make $350?

OpenStudy (anonymous):

Add them up

OpenStudy (anonymous):

a. b = 40 - p 8b + 4p = 40 b. b = 40 - p 42b + 52p = 350 c. 8b = 40 - 4p 42b - 52p = 350 d. 8b = 40 - 4p 42b + 52p = 350 @lainey06 i dont know which one. pls help

OpenStudy (anonymous):

wouldnt you start from B then P? @dxgx3

OpenStudy (anonymous):

@Lainey06 I have no clue

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@Hero @radar

OpenStudy (beccaboo333):

Do you have any idea how to do this?

OpenStudy (anonymous):

@beccaboo333 Not at all.

OpenStudy (beccaboo333):

@ganeshie8 Can help you with this. I have an idea on what to do, but I'm not very good at explaining :(

OpenStudy (anonymous):

@ganeshie8 Pleasee help! & thank you @beccaboo333 :)

OpenStudy (anonymous):

how about this one @beccaboo333

OpenStudy (anonymous):

@ganeshie8